# Divergences induced by dual subtractive and divisive normalizations of exponential families and their convex deformations

### 20 Dec 2023 v1

@@ 5.2 Dually flat spaces
Definition 1 (Legendre type function [28])
@@
When $F(\theta)$ is chosen as the cumulant function of an exponential family, the divergence $D(p_{\theta_{1}}:p_{\theta_{2}})$ can be interpreted as a statistical divergence between probability densities: Namely, the reverse Kullback-Leibler divergence.
Proof:
The proof of the statistical divergence reconstruction was given in [18] and recalled here for sake of completeness. First, we prove that $F^{*}(\eta)=E_{p_{\theta}}[\log p_{\theta}]=-H(p_{\theta})$:
@@
Notice that when $k(x)\not=0$, we have an additional term: $H(p_{\theta})=-F^{*}(\eta)-E_{p_{\theta}}[k(x)]$, see [24].
@@
Thus the canonical divergence induced by the cumulant function amounts to a reverse Kullback-Leibler divergence $D_{\mathrm{KL}}^{*}$ between probability densities. $\square$
@@
  • $F(\theta)$ is the cumulant function,
@@
Thus in general there are many more dually flat spaces with corresponding divergences and statistical divergences than the usually considered exponential family manifold [29, 19] induced by the cumulant function. It is interesting to consider their use in information sciences.
6 Conclusion and discussion
For the machine learning practioner, it is well-known that the Kullback-Leibler divergence (KLD) between two probability densities $p_{\theta_{1}}$ and $p_{\theta_{2}}$ of an exponential family with cumulant function $F$ amounts to a reverse Bregman divergence [4] induced by $F$ or equivalently to a reverse Fenchel-Young divergence [2]:
where $\eta=\nabla F(\theta)$ is the dual moment/expectation parameter.
In this note, we showed that the KLD extended to positive densities ${\tilde{p}}_{\theta_{1}}$ and ${\tilde{p}}_{\theta_{2}}$ of an exponential family with convex partition function $Z(\theta)$ amounts to a reverse Bregman divergence induced by $Z$ or equivalently to a reverse Fenchel-Young divergence:
where $\tilde{\eta}=\nabla Z(\theta)$.
More generally, we showed that scaled $\alpha$-skewed Jensen divergences induced by the cumulant and partition functions between natural parameters coincide with the scaled $\alpha$-skewed Bhattacharyya distances between probability densities and the $\alpha$-divergences between unnormalized densities, respectively.
Since the log-convex partition function is also convex, we generalized the principle of building pairs of convex generators using comparative convexity with respect to a pair of quasi-arithmetic means, and discussed about the induced dually flat spaces and divergences.
Finally, a class of statistical divergences called projective divergences are invariant under rescaling: For example, the $\gamma$-divergence [12] $D_{\gamma}$ is such that $D_{\gamma}(p:q)=D_{\gamma}({\tilde{p}}:{\tilde{q}})$, and the $\gamma$-divergence tends to the KLD when $\gamma\rightarrow 0$. Other projective divergences include the Cauchy-Schwarz divergence [27] and the family of Hölder divergences [26].

### 18 Jan 2024 v2

@@ 5.2 Dually flat spaces
Definition 2 (Legendre type function [35]).
@@
When $F(\theta)$ is chosen as the cumulant function of an exponential family, the Bregman divergence $B_{F}({\theta_{1}}:{\theta_{2}})$ can be interpreted as a statistical divergence between corresponding probability densities. Namely, the Bregman divergenec amounts to the reverse Kullback-Leibler divergence: $B_{F}({\theta_{1}}:{\theta_{2}})=D_{\mathrm{KL}}^{*}(p_{\theta_{1}}:p_{\theta_{2}})$ (see Appendix B).
Thus in general there are many more dually flat spaces with corresponding divergences and statistical divergences than the usually considered exponential family manifold [36] induced by the cumulant function. It is interesting to consider their use in information sciences.
6 Conclusion and discussion
For the machine learning practioner, it is well-known that the Kullback-Leibler divergence (KLD) between two probability densities $p_{\theta_{1}}$ and $p_{\theta_{2}}$ of an exponential family with cumulant function $F$ (free energy) amounts to a reverse Bregman divergence [5] induced by $F$ or equivalently to a reverse Fenchel-Young divergence [3]:
where $\eta=\nabla F(\theta)$ is the dual moment or expectation parameter.
In this paper, we showed that the KLD extended to positive unnormalized densities ${\tilde{p}}_{\theta_{1}}$ and ${\tilde{p}}_{\theta_{2}}$ of an exponential family with convex partition function $Z(\theta)$ (Laplace transform) amounts to a reverse Bregman divergence induced by $Z$ or equivalently to a reverse Fenchel-Young divergence:
where $\tilde{\eta}=\nabla Z(\theta)$.
More generally, we showed that the scaled $\alpha$-skewed Jensen divergences induced by the cumulant and partition functions between natural parameters coincide with the scaled $\alpha$-skewed Bhattacharyya distances between probability densities and the $\alpha$-divergences between unnormalized densities, respectively:
$D_{B,\alpha}^{s}(p_{\theta_{1}}:p_{\theta_{2}})$ $=$ $J_{F,\alpha}^{s}(\theta_{1}:\theta_{2}),$
$D_{\alpha}({\tilde{p}}_{\theta_{1}}:{\tilde{p}}_{\theta_{2}})$ $=$ $J^{s}_{Z,\alpha}(\theta_{1}:\theta_{2}).$
We noticed that the partition functions $Z$ of exponential families are both convex and log-convex, and the corresponding cumulant functions are both convex and exponentially convex.
Figure 3 summarizes the relationships between statistical divergences between normalized and unnormalized densities of an exponential family and corresponding divergences between their natural parameters. Notice that Brekelmans and Nielsen [10] considered deformed uni-order likelihood ratio exponential families (LREFs) for annealing paths, and obtained an identity between the $\alpha$-divergences between unnormalized densities and Bregman divergences induced by multiplicatively scaled partition functions.
Figure 3: Statistical divergences between normalized $p_{\theta}$ and unnormalized ${\tilde{p}}_{\theta}$ densities of an exponential family $\mathcal{E}$ with corresponding divergences between their natural parameters. Without loss of generality, we consider a natural exponential family (i.e., $t(x)=x$ and $k(x)=0$) with cumulant function $F$ and partition function $Z$. $J_{F}$ and $B_{F}$ denote the Jensen and Bregman divergences induced by generator $F$, respectively. Statistical divergences $D_{R,\alpha}$ and $D_{B,\alpha}$ denote the Rényi $\alpha$-divergences and the skewed $\alpha$-Bhattacharyya distances, respectively. The superscript “s” indicates a rescaling by the multiplicative factor $\frac{1}{\alpha(1-\alpha)}$, and the superscript “*” denote the reverse divergence obtained by swapping the parameter order.
Since the log-convex partition function is also convex, we generalized the principle of building pairs of convex generators using comparative convexity with respect to a pair of quasi-arithmetic means, and discussed about the induced dually flat spaces and divergences. In particular, by considering the convexity-preserving deformations obtained by power mean generators, we show how to obtain a family of convex generators and dually flat spaces. Notice that some parametric families of Bregman divergences like the $\alpha$-divergences [2] or the $\beta$-divergences [20] yield smooth families of dually flat spaces.
Concave deformations have also been recently studied in [22]: The authors introduce the $\log_{\phi}$-concavity induced by a positive continuous function $\phi$ generating a deformed logarithm $\log_{\phi}$ as the $(A,\log_{\phi})$-comparative concavity (Definition 1.2 in[22]), and the weaker notion of $F$-concavity which corresponds to the $(A,F)$-concavity (Definition 2.1 in[22], requiring strictly increasing functions $F$). Our deformation framework $Z=\tau^{-1}\circ F\circ\rho$ is more general since double-sided: We deform both the function $F$ by $F_{\tau}=\tau^{-1}\circ F$ and its argument $\theta$ by $\theta_{\rho}=\rho(\theta)$.
Exponentially concave functions are considered as generators of $L$-divergences in [40], and $\alpha$-exponentially concave functions $G$ such that $\exp(\alpha G)$ are concave for $\alpha>0$ generalize the $L$-divergences to $L_{\alpha}$-divergences which can be expressed equivalently using a generalization of the Fenchel-Young divergence based on the $c$-transforms [40]. When $\alpha<0$, exponentially convex functions are considered instead of exponentially concave functions. The information geometry induced by $L_{\alpha}$-divergences are dually projectively flat with constant curvature, and reciprocally a dually projectively flat structure with constant curvature induces (locally) a canonical $L_{-\alpha}$-divergence. Wong and Zhang [44] investigate a one-parameter deformation of convex duality called $\lambda$-duality by considering functions $f$ such that $\frac{1}{\lambda}(e^{\lambda f}-1)$ are convex for $\lambda\not=0$. They define the $\lambda$-conjugate transform as a particular case of the $c$-transform [40] and study the information geometry of the induced $\lambda$-logarithmic divergences. The $\lambda$-duality yields a generalization of exponential and mixture families to $\lambda$-exponential and $\lambda$-mixture families related to the Rényi divergence.
Finally, a class of statistical divergences called projective divergences are invariant under rescaling: For example, the $\gamma$-divergence [18] $D_{\gamma}$ is such that $D_{\gamma}(p:q)=D_{\gamma}({\tilde{p}}:{\tilde{q}})$, and the $\gamma$-divergence tends to the KLD when $\gamma\rightarrow 0$.
Appendix A Convexity of the cumulant functions of exponential families
Let us prove Proposition 1 and Proposition 2:
Proposition [7] The natural parameter space $\Theta$ of an exponential family is convex.
Proof.
Let $\Theta$ denote the natural parameter space:
since both $\int\exp({\left\langle\theta_{0},x\right\rangle})\mathrm{d}\mu(x)<\infty$ and $\int\exp({\left\langle\theta_{1},x\right\rangle})\mathrm{d}\mu(x)<\infty$ because $\theta_{0}$ and $\theta_{1}$ both belong to $\Theta$. Hence, we have shown that $\Theta$ is convex. ∎
Proposition [7] The cumulant function $F(\theta)$ is strictly convex and the partition function $Z(\theta)$ is positive and strictly log-convex.
Proof.
Eq. 32 can be rewritten as
Now, a function $Z(\theta)$ is strictly log-convex iff $\log Z(\theta)$ is strictly convex [29]. Since $Z(\theta)=\exp(F(\theta))$, we deduce that $Z(\theta)$ is positive and strictly log-convex since $\log Z(\theta)=F(\theta)$ is strictly convex. ∎
In Proposition 3, we prove that strictly log-convex functions are strictly convex. Let us prove a weaker version of this proposition using the second-order differentiability of convex functions:
Proposition A log-convex $C^{2}$ function $Z:\Theta\subset\mathbb{R}^{m}\rightarrow\mathbb{R}$ is convex.
Proof.
A $C^{2}$ function $f(x)$ is convex iff $f^{\prime\prime}(x)\geq 0$ but a $C^{2}$ function $f(x)$ with $f^{\prime\prime}(x)>0$ is strictly convex but not necessarily the converse: For example, $f(x)=x^{4}$ is strictly convex but $f^{\prime\prime}(x)=12x^{2}$ vanishes for $x=0$. We can prove the convexity of log-convex functions by considering successively the univariate and multivariate cases as follows:
  • In the univariate case $m=1$, consider the log-convex function $Z(\theta)=\exp({F(\theta)})$ for $F(\theta)$ a strictly convex function with $F^{\prime\prime}(\theta)\geq 0$. Then we have $Z^{\prime\prime}(\theta)=F^{\prime\prime}(\theta)e^{F(\theta)}+(F^{\prime}(\theta))^{2}e^{F(\theta)}\geq 0$.
  • In the multivariate case $m>1$, $C^{2}$ function $F(\theta)$ is strictly convex if its Hessian is positive semi-definite. Let $Z(\theta)$ be $C^{2}$ strictly log-convex, i.e., $F(\theta)=\log Z(\theta)$ is strictly convex. Since $Z(\theta)=e^{F(\theta)}$ for $\nabla^{2}F(\theta)\succeq 0$, we need to check that $\nabla^{2}Z(\theta)\succeq 0$. Let $\partial_{l}=\frac{\partial}{\partial\theta_{l}}$. Then we have
    or in matrix form
    Since (i) $e^{F(\theta)}>0$, (ii) $\nabla^{2}F(\theta)\succeq 0$ and (iii) $\nabla F(\theta)(\nabla F(\theta))^{\top}\succeq 0$ (since a matrix outer-product is positive semi-definite: $\theta^{\top}\nabla F(\theta)(\nabla F(\theta))^{\top}\theta=(\theta^{\top}\nabla F(\theta))^{2}\geq 0$), we deduce that $\nabla^{2}Z(\theta)$ is positive semi-definite. Hence, $Z(\theta)$ is convex.
Appendix B Statistical divergence recovered from the Bregman divergence induced by cumulant functions
Proposition The Bregman divergence $B_{F}({\theta_{1}}:{\theta_{2}})$ induced by the cumulant function $F(\theta)$ of an exponential family can be interpreted as a statistical divergence between corresponding probability densities $p_{\theta_{1}}$ and $p_{\theta_{2}}$: $B_{F}({\theta_{1}}:{\theta_{2}})=D_{\mathrm{KL}}^{*}(p_{\theta_{1}}:p_{\theta_{2}})$.
Proof.
First, we prove that $F^{*}(\eta)=E_{p_{\theta}}[\log p_{\theta}]=-H(p_{\theta})$:
@@ Appendix B Statistical divergence recovered from the Bregman divergence induced by cumulant functions
Notice that when $k(x)\not=0$, we have an additional term: $H(p_{\theta})=-F^{*}(\eta)-E_{p_{\theta}}[k(x)]$.
@@
Thus the canonical divergence induced by the cumulant function amounts to a reverse Kullback-Leibler divergence $D_{\mathrm{KL}}^{*}$ between probability densities. ∎
@@
  • $F(\theta)$ is the cumulant function (also called free energy in thermodynamics),